OpenStudy (anonymous):

Medal Simplify -3/5 - 9/10 - (-4/5)

OpenStudy (anonymous):

First, find a common denominator for the fractions.

OpenStudy (anonymous):

how would you find that?

OpenStudy (anonymous):

Well, we can always multiply a number by 1. But if we choose a good way to write the number 1, we can change what the fraction looks like without changing its value (this will make sense in a minute). Let's look at the first two. The denominator of the first fraction is 5. The denominator of the second fraction is 10. What do we need to multiply 5 by to get 10?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Good. So we'll do that. But we can't just multiply the bottom by 2, we also have to multiply the top by 2. That's because we want to multiply the entire fraction by 1, and 2/2 = 1 So we get \[\frac{3}{5} = \frac{3}{5}*\frac{2}{2} = \frac{3*2}{5*2} = \frac{6}{10}\] The third fraction has a denominator of 5. Can you use this approach to give it a denominator of 10?

OpenStudy (anonymous):

where do i start

OpenStudy (anonymous):

Exactly what we did with the first fraction.

OpenStudy (anonymous):

multiply by 2

OpenStudy (anonymous):

top and bottom, yes.

OpenStudy (anonymous):

Do what Vandreigan showed to all values that are not over 10

OpenStudy (anonymous):

6/10 * 2 = 12/20

OpenStudy (anonymous):

We're looking at 4/5. We need that to have a denominator (number on the bottom) of 10.

OpenStudy (anonymous):

* by 2 and the denominator will be 10

OpenStudy (anonymous):

Good, but we'll also need to multiply the top by 2. What will be the resulting fraction?

OpenStudy (anonymous):

8/10

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Excellent. Now we can rewrite the starting question. -6/10 - 9/10 - (-8/10) = ?

OpenStudy (anonymous):

There are no exponents or anything like that, so the two negative signs in front of the 8/10 will cancel. Then we just add/subtract like normal.

OpenStudy (anonymous):

-23/10

OpenStudy (anonymous):

Not quite -6/10 - 9/10 -(-8/10) = -6/10 - 9/10 + 8/10

OpenStudy (anonymous):

-7/10

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

That's it :)

OpenStudy (anonymous):

thank you :)